OpenStudy (anonymous):

How do you simplify the square root of 162x^5y^6??

OpenStudy (btaylor):

factor out 162 x^5 y^6. If a factor appears twice, put it outside the radical. Then condense it back together again. 162 x^5 y^6 = (2)(3)(3)(x)(x)(x)(x)(x)(y)(y)(y)(y)(y)(y). So you get 3x^2 y^3 outside the radical, and inside you have 2x. So the final answer: \[3x^{2}y^{3}\sqrt{2x}\]

OpenStudy (anonymous):

Thanks soo much!!